OpenStudy (anonymous):

find the trigonometric form of 4i.

OpenStudy (mertsj):

0+4i

OpenStudy (mertsj):

\[4\cos \frac{\pi}{2}+4\sin \frac{\pi}{2}\]

OpenStudy (anonymous):

could you explain, how though?

OpenStudy (mertsj):